iosangel:

At a zoo, the lion pen has a ring-shaped sidewalk around it. The outer edge of the sidewalk is a circle with a radius of 11 m. The inner edge of the sidewalk is a circle with a radius of 9 m. (A)Write and simplify an expression for the exact area of the sidewalk. (B)Find the approximate area of the sidewalk. Use 3.14 to approximate π. I don't want the answer, I just want to know how to get to the answer.....

iosangel:

Vocaloid:

A) the area of the sidewalk is the area of the big circle - area of small circle, so you would use the formula pi(r^2) to come up with a formula for each of the circles areas b) you would just evaluate/calculate whatever you wrote for part A), substituting 3.14 for pi lmk if you want to check your answer or if you're still stuck

iosangel:

A) This could be the expression pi(r^2)-pi(r^2) 11^2-9^2=40pi B) 11^2-9^2=40*3.14=125.6m^2 Is this correct?

Vocaloid:

yeah that's good, just be sure to write (pi)11^2 - (pi)9^2 = 40pi instead of just 11^2-9^2=40pi